Yoko deposits $2000 into an account that pays simple interest at a rate of 5% per year. how much interest will she be paid in the first 5 years?
Answer by ikleyn(52776)

.

The simple interest is


    I = P*r*t = 2000*0.05*5 = 500 dollars in the first 5 years.

Solved.